OpenStudy (adalrymple):

Horizontal number line with nine markings, Center mark is marked ten with a bold dot. Line to the right of ten is bolded. Right and left ends of line are arrows pointing outward. Select one: A. two fifths times p is greater than or equal to four B. negative two fifths times p is less than or equal to four C. negative two fifths times p is greater than or equal to four D. two fifths times p is less than or equal to four

OpenStudy (adalrymple):

if you want a clearer picture

OpenStudy (jdoe0001):

hmm what is that line saying anyway? it has a solid ball, then goes from 10 to the right what does that mean?

OpenStudy (adalrymple):

It means its greater than or equal to @jdoe0001

OpenStudy (jdoe0001):

ok let's see the given choices, to see which one might be let's see C \(\bf -\cfrac{2}{5}p\ge 4\implies -2p\ge5\cdot 4\implies -2p\ge 20 \\ \quad \\ -\cfrac{1}{2}\cdot -2p\ge-\cfrac{1}{2}\cdot 20\implies p{\color{blue}{ \le}}-10\) so.... that says, P is less than or equals to -10 so, that one is not it so.. check the others, BD and A see if you get \(\bf p \ge 10\)

OpenStudy (anonymous):

you have no choice, you have to solve each one and see which gives \[x\geq 10\]

OpenStudy (anonymous):

you do this by multiplying \(4\) by \(\frac{5}{2}\) for the first one